CH2Cl2 + HCI biphenyl major organic product biphenyl + t-BuCI + Al (foil) — > 4,4%-di-tert-butylbiphenyl MW/ Conc 26.98 volumfill in every blank in the table and show your work how you get the answer for each.
Answer #1

Biphenyl: Mass = 0.169 Molar mass: 154.2 g/mol Moles of biphenyl = mass /molar mass = 0.16g/154.2 g/mol = 1.038*10^-3 mol = 1biphenyi # -Buci + Alfoil) - > 4,4-di-tert-butylbiphenyi 26.98 MW/ Conc volume density 0.35 ml solid 0.851gm/L 0.298g Theo ma